OpenStudy (anonymous):

The sum of 3 consecutive odd integers is -21. Find the "middle" number. A. -9 B. -3 C. -7 D. -5

OpenStudy (raden):

let the numbers are x, x+2, and x+4 the expression will satisfied : x + x + 2 + x + 4 = -21 3x + 6 = -21 3x = -21 - 6 3x = -27 x = -27/3 x = -9 so, they are -9, -9+2, -9+4 or -9, -7, -5 as the middle is -7